# SDPose: Exploiting Diffusion Priors for Out-of-Domain and Robust Pose Estimation

### v1

@@ Abstract
Pre-trained diffusion models provide rich multi-scale latent features and are emerging as powerful vision backbones. While recent works such as Marigold (Ke et al., 2024) and Lotus (He et al., 2024) adapt diffusion priors for dense prediction with strong cross-domain generalization, their potential for structured outputs (e.g., human pose estimation) remains underexplored. In this paper, we propose SDPose, a fine-tuning framework built upon Stable Diffusion to fully exploit pre-trained diffusion priors for human pose estimation. First, rather than modifying cross-attention modules or introducing learnable embeddings, we directly predict keypoint heatmaps in the SD U-Net’s image latent space to preserve the original generative priors. Second, we map these latent features into keypoint heatmaps through a lightweight convolutional pose head, which avoids disrupting the pre-trained backbone. Finally, to prevent overfitting and enhance out-of-distribution robustness, we incorporate an auxiliary RGB reconstruction branch that preserves domain-transferable generative semantics. To evaluate robustness under domain shift, we further construct COCO-OOD, a style-transferred variant of COCO with preserved annotations. With just one-fifth of the training schedule used by Sapiens on COCO, SDPose attains parity with Sapiens-1B/2B on the COCO validation set and establishes a new state of the art on the cross-domain benchmarks HumanArt and COCO-OOD. Furthermore, we showcase SDPose as a zero-shot pose annotator for downstream controllable generation tasks, including ControlNet-based image synthesis and video generation, where it delivers qualitatively superior pose guidance.
@@ 1 Introduction
To bridge this gap and systematically study how to leverage the rich latent features for robust, cross-domain pose estimation, we propose SDPose, a fine-tuning framework with three key components: (i) Latent-space preservation. We operate entirely in the SD U-Net’s image latent space without modifying cross-attention modules or adding learnable embeddings, thus preserving pretrained visual semantics and feature geometry. (ii) Lightweight pose decoder head. We employ one deconvolution layer (kernel size 4) followed by two $1\times 1$ convolutions to map intermediate U-Net features directly into keypoint heatmaps, minimizing disruption to the pretrained backbone. (iii) RGB reconstruction regularization. We add an auxiliary reconstruction branch that regularizes fine-tuning, helping to maintain domain-transferable generative semantics and improve out-of-distribution robustness. Furthermore, to systematically evaluate robustness under domain shift, we introduce COCO-OOD, a style-transferred extension of COCO annotations featuring oil painting domains, which bridges the gap in benchmarking for generalization robustness evaluation.
@@ 4.1 Leveraging the Multi-scale Latent Features for Pose Estimation
In this paper, we directly leverage the multi-scale latent features of SD U-Net for the pose estimation task. The input image is encoded by the frozen SD-VAE encoder and then fed into the SD U-Net, from which we extract multi-scale features at the upsampling stage. These features serve as robust representations for downstream applications. The choice of feature map depends on the keypoint granularity: for the 17-keypoint setting, we use the last-layer upsampling features, while for the 133-keypoint whole-body setting, we use the second-to-last upsampling features, as validated in Sec. 5.4.2.
@@ 4.2 The U-Net Convolutional Output Layer Forms an Information Bottleneck
Stable Diffusion’s U-Net outputs a 4-channel latent $z\in\mathbb{R}^{4\times h\times w}$ for the VAE through a single convolutional layer. In contrast, pose estimation requires $K$-channel heatmaps $H\in\mathbb{R}^{K\times H^{\prime}\times W^{\prime}}$ with $K\gg 4$, making the 4-channel latent a severe information bottleneck. To address this, we replace the original 4-channel head with a lightweight heatmap decoder (Xiao et al., 2018). The decoder consists of a deconvolution layer for upsampling, followed by two $1\times 1$ convolutions that output $K$-channel heatmaps (Fig. 2). This modification removes the bottleneck and shortens the supervision path to keypoints.
@@ 5.1 Experiment Settings
Training Datasets. We train two variants, SDPose 17-keypoints and SDPose 133-keypoints, on MS COCO (Lin et al., 2014) and COCO-WholeBody (Jin et al., 2020), respectively. All images are processed using standard top-down augmentations, with the input resolution set to 1024 $\times$ 768. Further details are provided in the supplementary materials.
Validation Datasets and Metrics. (1) For the 17-keypoint variant, we evaluate SDPose on MS COCO (Lin et al., 2014) for real-world images, and on HumanArt (Ju et al., 2023) and COCO-OOD for cross-domain benchmarks. (2) For the 133-keypoint variant, we evaluate SDPose on COCO-WholeBody (Jin et al., 2020) and the extended COCO-OOD. Further details of the evaluation datasets and metrics are provided in the supplementary materials.
Figure 4: COCO-OOD visualization. COCO images are stylized into a Monet-like oil painting domain.
COCO-OOD (style-transferred COCO). To complement HumanArt and enable OOD evaluation with matched content/labels, we translate all COCO val images into Monet-style paintings using the official CycleGAN framework (Zhu et al., 2017) (monet2photo), preserving the original human annotations (boxes, keypoints). This yields an OOD split with unchanged scene geometry but shifted texture, color, and brush patterns. The domain shift is further quantified by a Fréchet Inception Distance (FID) (Heusel et al., 2017) of 46.23 between COCO-OOD and the original COCO validation images, which is notably larger than the 32.59 FID measured between COCO validation and HumanArt.
Please refer to the supplementary materials for additional details.
@@ 5.2 Quantitative and Qualitative Comparison on Real-World Scenes
Table 1: Quantitative comparison on the COCO validation set. We compare methods with parameter sizes comparable to the SDPose pipeline.
Model
Pre-trained
Backbone
Parameters Input size
Train
dataset
Train
epochs
AP AR
Sapiens-1B (Khirodkar et al., 2024) Sapiens ViT 1.169B 1024 $\times$ 768 COCO 210 82.1 85.9
Sapiens–2B (Khirodkar et al., 2024) Sapiens ViT 2.163B 1024 $\times$ 768 COCO 210 82.2 86.0
GenLoc (Wang et al., 2025a) Stable Diffusion-v1.5 0.95B N/A COCO 14 77.6 80.7
ViTPose ++-H  (Xu et al., 2023) ViTAE 0.63B 256 $\times$ 192 Mixture* 210 79.4 N/A
SDPose (Ours) Stable Diffusion-v1.5 0.95B 1024 $\times$ 768 COCO 40 81.2 85.3
SDPose (Ours) Stable Diffusion-v2 0.95B 1024 $\times$ 768 COCO 40 81.3 85.2
    Mixture includes COCO, AIC, MPII, AP10K, APT36K, and WholeBody datasets.
For the 17-keypoints variant, SDPose achieves 81.3 AP / 85.2 AR on the COCO validation set (Table 1) with only 40 training epochs using a 0.95B SD-v2 backbone. It matches the accuracy of Sapiens (82.1–82.2 AP) despite requiring 5× fewer epochs and a smaller backbone, and surpasses GenLoc (+3.7 AP, +4.5 AR). SDPose also outperforms ViTPose++ (+1.9 AP), which relies on multiple auxiliary datasets, while being trained solely on COCO. Figure 5 further illustrates robustness on real-world photos, where SDPose rivals Sapiens and corrects its failure cases (e.g., Sichuan opera eye keypoints). For the 133-keypoints variant, SDPose achieves competitive performance with Sapiens-1B on the COCO-WholeBody validation set. Further details are provided in the supplementary materials.
@@ 5.3 SDPose’s Strong OOD Robustness
Table 2: Quantitative comparison on OOD benchmarks HumanArt and COCO-OOD. Here, we report the methods with parameter sizes comparable to the SDPose pipeline. All models are trained on COCO with an input size of 1024 $\times$ 768; full HumanArt results for other methods are in the supplementary.
Model
Pre-trained
Backbone
Parameters
Train
epochs
HumanArt     COCO-OOD
    AP AR     AP AR
Sapiens-1B (Khirodkar et al., 2024) Sapiens ViT 1.169B 210     64.3 67.4     58.8 63.3
Sapiens-2B (Khirodkar et al., 2024) Sapiens ViT 2.163B 210     69.6 72.2     59.6 64.0
GenLoc (Wang et al., 2025a) Stable Diffusion-v1.5 0.95B 14     67.0 70.8     N/A N/A
SDPose (Ours) Stable Diffusion-v2 0.95B 40     71.2 73.9     63.5 68.2
In this section, we demonstrate the superior OOD robustness of SDPose using quantitative evaluation on HumanArt and our COCO-OOD benchmark. As shown in Table 2, SDPose achieves state-of-the-art results on HumanArt and COCO-OOD with fewer training epochs and a smaller parameter budget. On COCO-OOD WholeBody (Table 3), SDPose continues to demonstrate strong out-of-domain robustness. As shown in Fig. 1, SDPose achieves more accurate body pose estimation across diverse animation styles and humanoid robots compared with baseline models. Additional qualitative results on whole-body pose estimation in stylized paintings are provided in the supplementary materials.
Table 3: Quantitative comparison on the COCO-OOD Wholebody validation set. In this section, we report only the whole-body AP and AR; please refer to the supplementary materials for detailed results on individual body parts.
Model
Pre-trained
Backbone
Parameters Input size
Train
dataset
Train
epochs
Whole AP Whole AR
Sapiens-1B (Khirodkar et al., 2024) Sapiens ViT 1.169B 1024 $\times$ 768 COCO 210 38.7 46.8
Sapiens-2B (Khirodkar et al., 2024) Sapiens ViT 2.163B 1024 $\times$ 768 COCO 210 44.4 53.0
SDPose (Ours) Stable Diffusion-v2 0.95B 1024 $\times$ 768 COCO 42 46.6 54.8
@@ 5.4.1 Ablation on Auxiliary Reconstruction Task and Diffusion Priors
[…]
Method COCO     HumanArt     COCO-OOD
@@
SDPose 81.3 85.2     71.2 73.9     63.5 68.2
SDPose w/o RGB Recon. 80.8 (-0.5) 84.9 (-0.3)     69.8 (-1.4) 72.6 (-1.3)     62.5 (-1.0) 67.3 (-0.9)
SDPose w/o Diffusion Priors 74.9 (-6.4) 79.4 (-5.8)     53.8 (-17.4) 58.0 (-15.9)     52.7 (-10.8) 57.9 (-10.3)
@@ 5.4.2 Ablation on Multi-scale Features from the SD U-Net
Table 5: Ablation on U-Net feature selection for the pose head. All models are trained on COCO (40 epochs for 17-keypoints, 42 epochs for 133-keypoints).
Feature SDPose-17     SDPose-133
AP AR     AP AR
Last layer 81.3 85.2     70.5 77.5
Second-to-last layer 81.1 85.0     71.5 78.4
Prior work (Liu et al., 2023b; 2024; a; Wang et al., 2025b) suggests that penultimate features often transfer better than final ones. As shown in Table 5, SDPose-17 favors the last-layer feature, while SDPose-133 benefits from the second-to-last. This indicates that the optimal feature choice depends on keypoint granularity: coarse 17-keypoint settings are well served by the final representation, whereas whole-body localization gains from the richer semantics in the penultimate layer.
6 Zero-shot Downstream Applications
@@ 7 Conclusion
In this paper, we present SDPose, an SD-native fine-tuning strategy for human pose estimation. SDPose preserves the original U-Net with only lightweight task-specific components, adapting generative latent priors for keypoint prediction through an auxiliary RGB reconstruction branch and a heatmap decoder. We further introduce COCO-OOD, a style-transferred extension of COCO for evaluating robustness under domain shifts. With only 1/5 of Sapiens’ fine-tuning cost and a smaller backbone, SDPose matches its accuracy on COCO and achieves state-of-the-art results on COCO-OOD and HumanArt, demonstrating strong generalization for pose estimation.

### v2

@@ Abstract
Pre-trained diffusion models provide rich multi-scale latent features and are emerging as powerful vision backbones. While recent works such as Marigold (Ke et al., 2024) and Lotus (He et al., 2024) adapt diffusion priors for dense prediction with strong cross-domain generalization, their potential for structured outputs (e.g., human pose estimation) remains underexplored. In this paper, we propose SDPose, a fine-tuning framework built upon Stable Diffusion to fully exploit pre-trained diffusion priors for human pose estimation. First, rather than modifying cross-attention modules or introducing learnable embeddings, we directly predict keypoint heatmaps in the SD U-Net’s image latent space to preserve the original generative priors. Second, we map these latent features into keypoint heatmaps through a lightweight convolutional pose head, which avoids disrupting the pre-trained backbone. Finally, to prevent overfitting and enhance out-of-distribution robustness, we incorporate an auxiliary RGB reconstruction branch that preserves domain-transferable generative semantics. To evaluate robustness under domain shift, we further construct COCO-OOD, a style-transferred variant of COCO with preserved annotations. With just one-fifth of the training schedule used by Sapiens on COCO, SDPose attains parity with Sapiens-1B/2B on the COCO validation set and establishes a new state of the art on the cross-domain benchmarks HumanArt and COCO-OOD. Extensive ablations highlight the importance of diffusion priors, RGB reconstruction, and multi-scale SD U-Net features for cross-domain generalization, and t-SNE analyses further explain SD’s domain-invariant latent structure. We also show that SDPose serves as an effective zero-shot pose annotator for controllable image and video generation.
@@ 1 Introduction
To bridge this gap and investigate how SD’s rich latent representations can be effectively leveraged for robust cross-domain pose estimation, our contributions are as follows:
(1) We propose SDPose, a fine-tuning framework with three key components: (i) Latent-space preservation. We operate entirely in the SD U-Net’s image latent space without modifying cross-attention modules or adding learnable embeddings, thus preserving pretrained visual semantics and feature geometry. (ii) Lightweight pose decoder head. We introduce a minimal decoder that maps SD U-Net features to keypoint heatmaps with only a shallow convolutional head, ensuring low overhead and minimal disturbance to the pretrained latent representations. (iii) RGB reconstruction regularization. We add an auxiliary reconstruction branch that regularizes fine-tuning, helping to maintain domain-transferable generative semantics and improve out-of-distribution robustness.
(2) To systematically evaluate robustness under domain shift, we introduce COCO-OOD, a style-transferred extension of COCO that includes oil-painting, ukiyo-e, and color sketch domains. This dataset fills an important gap in benchmarking generalization robustness.
(3) We conduct extensive ablation studies to understand how diffusion priors and our RGB reconstruction branch contribute to cross-domain generalization in pose estimation. We further compare multi-scale features from different upsampling blocks of the SD-UNet, identifying the feature level that yields the strongest robustness under artistic domain shifts. Finally, through a latent-space comparison with Sapiens using t-SNE visualizations, we observe that SD’s pretrained latent features naturally capture domain-invariant structures, which is highly beneficial for cross-domain perception tasks.
@@ 4.1 Leveraging the Multi-scale Latent Features for Pose Estimation
We directly leverage the multi-scale latent features of SD U-Net for the pose estimation task. The input image is encoded by the frozen SD-VAE encoder and then fed into the SD U-Net, from which we extract multi-scale features at the upsampling stage. These multi-scale features provide rich and robust visual representations for downstream keypoint prediction. The specific feature level used for each task configuration is discussed in Sec. 5.5.2.
@@ 4.2 The U-Net Convolutional Output Layer Forms an Information Bottleneck
Stable Diffusion’s U-Net outputs a 4-channel latent $z\in\mathbb{R}^{4\times h\times w}$ for the VAE through a single convolutional layer. In contrast, pose estimation requires $K$-channel heatmaps $H\in\mathbb{R}^{K\times H^{\prime}\times W^{\prime}}$ with $K\gg 4$, making the 4-channel latent a severe information bottleneck. To address this, we replace the original 4-channel head with a lightweight heatmap decoder (Xiao et al., 2018). The decoder consists of deconvolution layers for upsampling, followed by convolutions that output $K$-channel heatmaps (Fig. 2). This modification removes the bottleneck and shortens the supervision path to keypoints.
@@ 4.4 Inference
Table 1: Quantitative comparison across COCO, HumanArt, COCO-OOD Monet, and COCO-OOD Ukiyo-e. All models are trained on COCO. Full quantitative comparison across various models are in the supplementary.
Model Variant  Model
Pre-trained
Backbone
Params
COCO
HumanArt
COCO-OOD
Monet
COCO-OOD
Ukiyo-e
        AP AR  AP AR  AP AR  AP AR
Body  Sapiens-1B (Khirodkar et al., 2024)  Sapiens ViT    1.169B    82.1 85.9    64.3 67.4    58.8 63.3    61.5 66.2
  Sapiens-2B (Khirodkar et al., 2024)  Sapiens ViT    2.163B    82.2 86.0  69.6 72.2    59.6 64.0    62.3 66.8
  GenLoc (Wang et al., 2025a)  Stable Diffusion-v1.5    0.95B    77.6 80.7    67.0 70.8    N/A N/A    N/A N/A
  SDPose (Ours)  Stable Diffusion-v2    0.95B    81.3 85.2    71.2 73.9  64.3 68.9  66.0 70.7
Wholebody  Sapiens-1B (Khirodkar et al., 2024)  Sapiens ViT    1.169B    72.7 79.2    N/A N/A    38.7 46.8    40.5 49.4
  Sapiens-2B (Khirodkar et al., 2024)  Sapiens ViT    2.163B    74.4 81.0  N/A N/A    44.4 53.0    46.6 55.8
  SDPose (Ours)  Stable Diffusion-v2    0.95B    71.5 78.4    N/A N/A    46.6 54.8  47.7 56.4
@@ 5.1 Experiment Settings
Training Datasets. We train two variants, SDPose Body (17-keypoints) and SDPose Wholebody (133-keypoints), on MS COCO (Lin et al., 2014) and COCO-WholeBody (Jin et al., 2020), respectively. All images are processed using standard top-down augmentations, with the input resolution set to 1024 $\times$ 768. Further details are provided in the supplementary materials.
Validation Datasets and Metrics. (1) For the Body variant, we evaluate SDPose on MS COCO (Lin et al., 2014) for real-world images, and on HumanArt (Ju et al., 2023) and COCO-OOD for cross-domain benchmarks. (2) For the Wholebody variant, we evaluate SDPose on COCO-WholeBody (Jin et al., 2020) and the extended COCO-OOD. Further details of the evaluation datasets and metrics are provided in the supplementary materials.
COCO-OOD. To complement HumanArt and enable OOD evaluation with matched content and labels, we translate all COCO val images into three artistic domains: Monet-style paintings using the official StyTR2 framework (Deng et al., 2022), ukiyo-e style using the official CycleGAN implementation (Zhu et al., 2017), and color-sketch style using Nano Banana (Nano Banana, ). For all stylized images, we reuse the original COCO val annotations (bounding boxes and keypoints). Please refer to the supplementary materials for details.
Figure 5: Illustration of the COCO-OOD dataset.
@@ 5.2 Quantitative and Qualitative Comparison on Real-World Scenes
For the Body variant, SDPose achieves 81.3 AP / 85.2 AR on the COCO validation set (Table 1) with only 40 training epochs using a 0.95B SD-v2 backbone. It matches the accuracy of Sapiens (82.1–82.2 AP) despite requiring 5× fewer epochs and a smaller backbone, and surpasses GenLoc (+3.7 AP, +4.5 AR). Figure 4 further illustrates robustness on real-world photos, where SDPose rivals Sapiens and corrects its failure cases (e.g., Sichuan opera eye keypoints). For the Wholebody variant, SDPose achieves competitive performance with Sapiens-1B on the COCO-WholeBody validation set. Further details are provided in the supplementary materials.
@@ 5.3 SDPose’s Strong OOD Robustness
In this section, we demonstrate the superior OOD robustness of SDPose using quantitative evaluation on HumanArt and our COCO-OOD benchmark. As shown in Table 1, SDPose achieves state-of-the-art results on HumanArt and COCO-OOD with fewer training epochs and a smaller parameter budget. On COCO-OOD WholeBody, SDPose continues to demonstrate strong out-of-domain robustness. As shown in Fig. 1, SDPose achieves more accurate body pose estimation across diverse animation styles and humanoid robots compared with baseline models. Additional qualitative results on whole-body pose estimation in stylized paintings are provided in the supplementary materials.
5.4 Why Diffusion Priors Exhibit Strong OOD Robustness?
In this subsection, we investigate why the diffusion-based prior (Stable Diffusion v2) (Rombach et al., 2022) exhibits stronger cross-domain generalization than the human-centric Sapiens ViT backbone (Khirodkar et al., 2024). We compare latent feature distributions from the pretrained Sapiens-1B ViT and Stable Diffusion v2 U-Net, as well as their pose-finetuned counterparts, Sapiens-1B Pose and our SDPose model (finetuned on the last four upsampling blocks). We sample 300 person instances from the COCO and COCO-OOD validation sets. For each instance, we crop the person region using the ground-truth bounding box and collect four stylistic variants: the original COCO image plus its ukiyo-e, Monet-oil, and color-sketch versions from COCO-OOD, yielding 1200 crops in total. These images are passed through the corresponding backbones to extract latent features, on which we run t-SNE (Maaten & Hinton, 2008) and compute silhouette scores (Rousseeuw, 1987).
Fig. 6(a–e) show the t-SNE visualizations of the pretrained priors, and Fig. 6(f–j) show the pose-finetuned models. In Fig. 6(a), the pretrained Sapiens features form clear style-driven clusters (silhouette by style =0.3469=0.3469) with a negative silhouette by person-instance (0.1608-0.1608), indicating that its representation is dominated by artistic appearance rather than instance structure. In contrast, Stable Diffusion U-Net features in Fig. 6(b–e) exhibit much weaker style separation and gradually stronger person-instance coherence; deeper upsampling blocks yield increasingly person-consistent and style-invariant distributions, reflected by rising person-instance silhouettes (up to $\approx 0.22$) and near-zero style silhouettes. After pose finetuning, both Sapiens-Pose (Fig. 6(f)) and SDPose (Fig. 6(g–j)) produce visible person-instance clusters, but SDPose features are noticeably tighter and more focused, with the mid-level SDPose blocks achieving the highest person silhouettes ($\approx 0.45$$0.48$) while keeping style silhouettes close to zero. Overall, the SD U-Net (especially after SDPose finetuning) provides more instance-coherent and style-disentangled representations than Sapiens, which explains why SD-based features offer stronger cross-domain robustness under artistic style shifts.
@@ 5.5 Ablation Study
Figure 6: t-SNE visualization of features from Sapiens ViT, Stable Diffusion U-Net blocks, Sapiens Pose, and SDPose across four visual domains. Each point corresponds to an image sample; colors represent person instances and marker shapes denote artistic styles.
@@ 5.5.1 Ablation on Auxiliary Reconstruction Task and Diffusion Priors
Table 2: Ablation studies on diffusion priors, RGB reconstruction, and U-Net feature selection. All experiments are trained on COCO with 40 epochs (body-17 keypoints) or 42 epochs (wholebody-133 keypoints).
Model Variant  Ablation Setting  COCO  HumanArt  COCO-OOD Monet
@@
SDPose-Body  Full Model (Last Block)    81.3 85.2  71.2 73.9  63.5 68.2
  w/o RGB Reconstruction    80.8 (-0.5) 84.9 (-0.3)    69.8 (-1.4) 72.6 (-1.3)    62.5 (-1.0) 67.3 (-0.9)
  w/o Diffusion Priors    74.9 (-6.4) 79.4 (-5.8)    53.8 (-17.4) 58.0 (-15.9)    52.7 (-10.8) 57.9 (-10.3)
  Last Block (F1)    81.3 85.2  71.2 73.9  63.5 68.2
  Second-to-last Block (F2)    81.1 85.0    70.4 73.3    64.3 68.9
  Third-to-last Block (F3)    81.0 85.1    70.6 73.3    63.5 68.2
  Fourth-to-last Block (F4)    79.2 83.4    65.0 68.1    58.1 62.8
SDPose-Wholebody  Last Block (F1)    70.5 77.5    N/A N/A    44.7 53.0
  Second-to-last Block (F2)    71.5 78.4  N/A N/A    46.6 54.8
  Third-to-last Block (F3)    70.4 77.6    N/A N/A    45.5 53.8
  Fourth-to-last Block (F4)    64.6 72.1    N/A N/A    37.1 45.4
@@ 5.5.2 Ablation on Multi-scale Features from the SD U-Net
Prior work (Liu et al., 2023b; 2024; a; Wang et al., 2025b) suggests that penultimate features often transfer better than final ones. As shown in Table 2, we evaluate SD U-Net upsampling features from the last four blocks. On COCO (17-keypoint), the last-block (F1) and penultimate (F2) features perform very similarly, and F1 is slightly better on both COCO and HumanArt, which contain a mixture of natural and artistic content. In contrast, on COCO-OOD Monet, where the domain shift is purely stylistic, F2 achieves the best performance, suggesting that it captures style-invariant cues more effectively than F1. For the 133-keypoint WholeBody setting, F2 consistently outperforms all other feature levels on both COCO-WholeBody and COCO-OOD, indicating that it offers a better balance between semantic robustness and spatial detail in this more fine-grained regime. Together with the latent-space analysis in Section 5.4, these results show that deeper SD features (F3 and F4) encode even stronger instance-consistent, style-invariant semantics but operate at much coarser spatial resolutions (H/16 and H/32), which removes the geometric details needed for accurate keypoint localization. In practice, we therefore use F1 as the default feature level for the SDPose Body model (favoring slightly better in-domain and mixed-domain performance) and F2 as the default feature level for the SDPose WholeBody model (favoring consistently stronger performance under both in-domain and out-of-domain settings).
6 Downstream Applications
@@ 7 Conclusion
In this paper, we present SDPose, an SD-native fine-tuning framework for human pose estimation. SDPose preserves the original U-Net with only lightweight task-specific components and adapts diffusion latent priors for keypoint prediction through an RGB reconstruction branch and a heatmap decoder. We further introduce COCO-OOD, a style-transferred extension of COCO for evaluating robustness under domain shifts. With only one-fifth of the fine-tuning cost of Sapiens and a smaller backbone, SDPose matches its in-domain accuracy on COCO and achieves state-of-the-art results on COCO-OOD and HumanArt. Our ablations and latent-space analyses show that diffusion priors and multi-scale SD features naturally encode domain-invariant structure, enabling strong generalization in both pose estimation and zero-shot pose annotation for controllable generation.